OpenStudy (anonymous):

simplify the expression: (4x^3y^4z^5)(5x^4y^5z^3)

OpenStudy (anonymous):

20x^7y^9z^8

OpenStudy (anonymous):

\[\large (4x^3y^4z^5)(5x^4y^5z^3) \implies (4 \times 5)(x^3 \times x^4)(y^4 \times y^5) (z^5 \times z^3)\]

OpenStudy (anonymous):

And remember in Multiplication bases are same then we add their exponents.. \[\large \color{green}{x^a \times x^b = x^{a+b}}\]

OpenStudy (goformit100):

@ghass1978 you can't just directly answer the question

OpenStudy (anonymous):

20 z^8 y^9 x^7

OpenStudy (anonymous):

multiply x^3 and 4 then x^3 by 1 answer is x^3 out then 4*x^3 and multply y^4 and 4x^3 and multiply y^4 and x^3 y^4 and x^3 then 4* x^3 * x^4 then z^5 * 4 y^4 * x^3 now we have 4* x^3 * y^4 * z^5 Multiply the exponent of y by y^5 Multiply the exponent of x by x^4 then 5 * x^4 you get 5*x^4*y^5 Multiply 4z^5*y^4*x^3 and 5*Z3*y^5*x^4 simplify and you will get z^8 y^9 x^7 4*5 is 20 so 20 z^8 y^9 x^7 Ans